# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### v4

@@ Abstract.
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past decade. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. MToP incorporates over 50 MTEAs, more than 200 MTO problem cases with real-world applications, and over 20 performance metrics. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 50 popular single-task evolutionary algorithms to address MTO problems. MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm
@@ 1. Introduction
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang et al., 2020; Tian et al., 2021b), multi-objective optimization (Zhang and Li, 2007; Deb and Jain, 2014), and combinatorial optimization (Feng et al., 2021c, a). In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei et al., 2022). A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
(1) $\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,\bm{F}_{k}(\bm{x}_{K})\Big{]},\\ \textit{s.t. }\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$
where $\Omega_{k}$ and $\bm{F}_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective optimization task, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task. For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta et al., 2016) as follows:
@@
In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. EMT has found successful applications in various domains, including engineering scheduling (Feng et al., 2021c; Gupta et al., 2022), nonlinear equation systems (Li et al., 2024e), feature selection in machine learning (Chen et al., 2022), anomaly detection (Wang et al., 2022b), point cloud registration (Wu et al., 2024a, b), and reinforcement learning (Zhang et al., 2023).
@@
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan et al., 2023; Gupta et al., 2018). Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan et al., 2021; Li et al., 2024f). The first attempt of EMT can be traced back to the multifactorial EA (Gupta et al., 2016), which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali et al., 2020; Zhou et al., 2021b), or a multi-population framework allocating separate populations for each task (Li et al., 2020; Wang et al., 2022a). Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta et al., 2016; Zhou et al., 2021b), autoencoding (Feng et al., 2019b; Zhou et al., 2021a), affine transformation (Xue et al., 2022; Lin et al., 2024; Gao et al., 2022), and adversarial generative models (Liang et al., 2023) have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali et al., 2020; Jiang et al., 2023a; Wu et al., 2024c), knowledge selection (Wang et al., 2022a; Liang et al., 2022c), and historical feedback (Li et al., 2022c; Lin et al., 2021; Tang et al., 2021b), have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs. Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field. Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field. In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization. For instance, PlatEMO (Tian et al., 2017) is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems. EvoX (Huang et al., 2024) is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training. EDOLAB (Peng et al., 2023) is an optimization tool specifically for dynamic optimization.
However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks. Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics. Specifically, single-task metrics evaluate the performance of each task in the MTO problem, while multitask metrics assess the overall performance of the MTO problem. These bring uncertainty to the pattern of results display and analysis. These distinctive requirements present significant challenges for the implementation of MTEAs in existing tools and platforms. Moreover, EMT has been extended to the subfields of many-task optimization (Chen et al., 2020; Liaw and Ting, 2019), multi-objective MTO (Gupta et al., 2017; Gao et al., 2023), competitive MTO (Li et al., 2022c, 2023a), and constrained multi-task optimization (Li et al., 2022a). Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. MToP is designed to provide a comprehensive software platform for researchers to evaluate MTEAs on benchmark MTO problems and explore real-world applications. The reason for using MATLAB is that it is a widely used programming language in the field of EC. Moreover, MATLAB provides a user-friendly graphical interface and a powerful numerical computing environment, making it an ideal choice for developing MToP. More importantly, the existing MTEAs and MTO problems are mainly implemented in MATLAB, which can be directly integrated into MToP. The main contributions of MToP are summarized as follows:
@@
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 50 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types. Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 50 single-task EAs of diverse types. In terms of synthetic test problems, MToP incorporates over 200 benchmark MTO problems alongside several real-world applications. Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities.
@@
The rest of this paper is organized as follows. Section 2 elaborates on the architecture of MToP. Section 3 describes approaches for performing experiments with MToP. Section 4 introduces how to extend the contents of MToP. Finally, the conclusion and future work are given in Section 5.
@@ 2. Architecture of MToP
MToP consists of key modules including the test, experiment, and data processing modules, which facilitate algorithm execution, visualization, and data analysis. These components allow for flexible experimentation through both GUI and command line interfaces. Following standardized code patterns, MToP integrates various algorithms and performance metrics.
@@ 2.1.1. Test Module
The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms. It facilitates running the selected algorithm on the chosen problem within MToP. The module includes visualization functions for both single-objective and multi-objective MTO problems. Users can select different options to display depending on the problem type, resulting in different visual outputs, as demonstrated below. In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig. 1 (a)-(b). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be visualized, as illustrated in Fig. 1 (c). Regarding multi-objective MTO problems, the Pareto front figures can be generated for multiple tasks, as depicted in Fig. 1 (d). Furthermore, the convergence behavior of metrics can be plotted after executing algorithms on problems, as exemplified in Fig. 1 (e).
In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior. Figure 2 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space. CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives. The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$. The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.4$. Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
(a) Gen=1
(b) Gen=100
(c) Gen=200
(d) Gen=500
(e) Gen=700
(f) Gen=1000
(g) Gen=1
(h) Gen=100
(i) Gen=200
(j) Gen=500
(k) Gen=700
(l) Gen=1000
Figure 2. Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations. The top row (a)-(f) shows the population in the objective space, while the bottom row (g)-(l) displays the population in the decision space.
In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=100 to 200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task. The schematics indicate that the population for the second task occupies a similar position in the decision space as the first task. However, these two tasks do not precisely overlap in the decision space. This observation suggests the presence of negative knowledge transfer, which could be contributing to this discrepancy. In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
Figure 3. The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
The workflow of API calls for executing an algorithm on a problem is depicted in Fig. 3. The process initiates with the <Start_Button> event in <MTO_GUI.m>, triggering the parameter setting API SetParameter() within the selected algorithm <MFEA.m>. Subsequently, it proceeds to the Run() function of the algorithm, which initially invokes the Initialization() API of the algorithm base class <Algorithm.m>, and then iterates through the function evaluation API Evaluation(). During the initialization phase, the parameter settings of the problem <CEC17_MTP1.m> to be solved are obtained via the problem base class <Problem.m>. Similarly, the Evaluation() function retrieves the specific optimization task object through the problem base class. Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function <Obj.m> via the CallMetric() API.
All functions within MToP that execute algorithms, including subsequent experiment modules, adhere to this procedure. Additionally, both algorithm and problem parameters can be directly modified through the GUI of MToP.

### v5

# MToP: A MATLAB Benchmarking Platform for Evolutionary Multitasking

@@ Abstract.
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past decade. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a need for a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source benchmarking platform, named MToP, for EMT. MToP incorporates over 50 MTEAs, more than 200 MTO problem cases with real-world applications, and over 20 performance metrics. Based on these, we provide benchmarking recommendations tailored for different MTO scenarios. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 50 popular single-task evolutionary algorithms to address MTO problems. Notably, we release extensive pre-run experimental data on benchmark suites to enhance reproducibility and reduce computational overhead for researchers. MToP features a user-friendly graphical interface, facilitating results analysis, data export, and schematic visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at: https://github.com/intLyc/MTO-Platform
Evolutionary multitasking, benchmarking platform, multitask optimization problem, evolutionary algorithm
@@ 1. Introduction
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and ease of implementation. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang2020CORCO; Tian2021CCMO), multi-objective optimization (Zhang2007MOEA-D; Deb2014NSGA3), and combinatorial optimization (Feng2021VRPHTO; Feng2021EEMTA). In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei2021Review-MTO). A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ minimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
[…]
where $\Omega_{k}$ and $\bm{F}_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective multi-task optimization problem, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of Pareto sets for the $k$-th task. For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta2016MFEA) as follows:
@@
In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During objective function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. EMT has found successful applications in various domains, including engineering scheduling (Feng2021VRPHTO; Gupta2021Half), nonlinear equation systems (Li2024Evolutionary), feature selection in machine learning (Chen2021Evolutionary), anomaly detection (Wang2022AUC), point cloud registration (Wu2024Evolutionary; Wu2023Evolutionary), and reinforcement learning (Zhang2023Multitask).
@@
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan2023Knowledge; Gupta2018Insights). Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan2021ETO; Li2024MFEA-RL). The first attempt of EMT can be traced back to the multifactorial EA (Gupta2016MFEA), which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali2020MFEA2; Zhou2021MFEA-AKT), or a multi-population framework allocating separate populations for each task (Li2020MFMP; Li2023MTEA-SaO; Li2024MTES-KG). Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta2016MFEA; Zhou2021MFEA-AKT), auto-encoding (Feng2019EMEA; Zhou2021Learnable; Gu2025MTEA-PAE), affine transformation (Xue2022AT-MFEA; Lin2024Ensemble), and adversarial generative models (Liang2023EMT-GS) have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali2020MFEA2; Jiang2023BoKT), knowledge selection (Wang2022MTEA-AD; Liang2022EMaTO-MKT; Li2025MTDE-MKTA), and historical feedback (Li2022CompetitiveMTO; Lin2021EMT-ET; Li2023MTSRA), have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
To overcome these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. MToP is designed to provide a comprehensive software platform for researchers to evaluate MTEAs on benchmark MTO problems and explore real-world applications. The selection of MATLAB is based on several strategic factors. First, many foundational and seminal works in the EMT field, including the original MFEA (Gupta2016MFEA) and MFEA-II (Bali2020MFEA2), were developed and open-sourced in MATLAB. This established a strong precedent and a rich ecosystem of existing code for the community. Second, the prominence of other successful EC platforms in MATLAB, most notably PlatEMO (Tian2017PlatEMO), provides a high-quality open-source reference. This synergy allows for the adaptaion and reuse of well-vetted modules, such as those for multi-objective optimization, which is crucial for the multi-objective MTO subfield. Finally, MATLAB’s powerful numerical computing environment and its capabilities for building user-friendly graphical interfaces make it an ideal choice for a comprehensive platform. The main contributions of MToP are summarized as follows:
@@
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 50 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types. Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 50 single-task EAs of diverse types. In terms of synthetic test problems, MToP incorporates over 200 benchmark MTO problems alongside several real-world applications. Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities. We also release extensive pre-run experimental data to enhance reproducibility and reduce computational overhead for researchers.
@@
The rest of this paper is organized as follows. Section 2 reviews related software platforms in EC and justifies the need for MToP. Section 3 elaborates on the architecture of MToP. Section 4 provides guidelines for using MToP. Section 5 presents the experimental validation of MToP. Finally, Section 6 gives discussions and outlook of MToP.
2. Related Work
As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs. Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field. Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field. In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization:
  • PlatEMO (Tian2017PlatEMO): A comprehensive MATLAB platform tailored for evolutionary multi-objective optimization.
  • EDOLAB (Peng2023EDOLAB): A MATLAB-based platform focusing on evolutionary dynamic optimization.
  • IOHprofiler (Doerr2018IOHprofiler): A modular framework, including IOHexperimenter (deNobel2024IOHexperimenter) and IOHanalyzer (Wang2022IOHanalyzer), for the detailed benchmarking and analysis of iterative optimization heuristics, with a primary focus on single-objective and multi-objective problems.
  • EvoX (Huang2024EvoX): A distributed GPU-accelerated library focused on scalability and expediting complex optimization and reinforcement learning tasks.
  • PyPop7 (Duan2024PyPop7): A pure-Python library for population-based single-objective black-box optimization, particularly for large-scale optimization.
  • MetaBox (Ma2023MetaBox): A Python-based platform for meta-black-box optimization, which uses meta-learning to design optimizers, with its latest version using MTO as a training scenario.
Table 1. Comparison of core focus areas among popular evolutionary computation platforms.
Platform Language Core Focus MTO Support
PlatEMO (Tian2017PlatEMO) MATLAB Evolutionary Multi-objective Optimization Partial (v4)
EDOLAB (Peng2023EDOLAB) MATLAB Evolutionary Dynamic Optimization No
IOHprofiler (Doerr2018IOHprofiler) C++/Python/R Benchmarking Iterative Optimization Heuristics No
EvoX (Huang2024EvoX) Python Distributed GPU-accelerated Optimization No
PyPop7 (Duan2024PyPop7) Python Large-scale Black-box Optimization No
MetaBox (Ma2023MetaBox) Python Meta-black-box Optimization via Meta-learning As a test scenario (v2)
MToP (This work) MATLAB Evolutionary Multitask Optimization Yes (Native)
A review of these tools, summarized in Table 1, justifies the development of a new, dedicated framework. At the time MToP development commenced [^1] MToP has been continuously updated since Sep. 21, 2021 at: https://github.com/intLyc/MTO-Platform, a platform with native, comprehensive support for MTO benchmarking was absent. The most prominent existing platform, PlatEMO (Tian2017PlatEMO), has a core architecture that is multi-objective optimization-centric, meticulously designed for solving single-task multi-objective optimization. This presents a fundamental architectural mismatch for EMT. Unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks, the implementation of inter-task solution space mapping, and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics, which brings uncertainty to the pattern of results display and analysis. These distinctive requirements present significant challenges for implementation in platforms not designed for them. Consequently, extending current framework like PlatEMO to natively handle these MTO-specific concepts is non-trivial and leads to significant extensibility limitations. This is particularly evident in the context of many-task optimization, where PlatEMO’s support for specialized multitask metrics is restricted. Although preliminary MTO support was added to PlatEMO (v4.0) [^2] PlatEMO v4.0 (released on Oct. 13, 2022) introduced preliminary MTO support., these additions are supplementary rather than architecturally native. The other platforms listed in Table 1 are similarly specialized for different, non-MTO-centric goals. Moreover, EMT has been extended to the subfields of many-task optimization (Chen2020MaTDE; Liaw2019SBO; Li2024TNG-NES), multi-objective MTO (Gupta2017MO-MFEA; Li2025MTEA-DCK), competitive MTO (Li2022CompetitiveMTO; Li2023MTSRA; Li2024CMO-MTO), and constrained multi-task optimization (Li2022CMTO-Benchmark; Zhang2024CEDA). This growing complexity further necessitates a dedicated, MTO-native framework. While a new framework was deemed necessary, MToP actively integrates and appropriately cites code modules from these established platforms where feasible, thereby reducing redundant development and focusing on its core MTO-specific contributions.
@@ 3. Architecture of MToP
In this section, we present the architecture of MToP, encompassing its functional modules, project structure, and code patterns.
@@ 3.1. Functional Modules
MToP is organized into three interconnected functional modules to support a full research workflow. Researchers can first use the Test Module for preliminary analysis and visualization of specific algorithms and problems. The Experiment Module facilitates large-scale, multi-run comparative experiments. Finally, the Data Process Module provides tools to manage the datasets generated by the Experiment Module, such as merging or splitting results for flexible post-processing. The following subsections detail the architectural role and capabilities of these modules. A practical guide on their specific operation and usage is provided in Section 4.
@@ 3.1.1. Test Module
The Test Module is designed to assist researchers in the qualitative analysis of MTO problems and algorithms. It facilitates executing a single run of a selected algorithm on a chosen problem to generate a suite of diagnostic visualizations. These tools help inspect both problem characteristics and algorithmic performance, as illustrated in Fig. 1. For single-objective MTO problems, users can depict function landscapes in one or two dimensions (Fig. 1 (a)-(b)) or visualize the feasible regions of constrained tasks (Fig. 1 (c)). For multi-objective MTO problems, the module can plot the true Pareto fronts with populations (Fig. 1 (d)).
In addition to these static plots, the module provides dynamic visualization utilities. Once a run is initiated, researchers can enable options such as Draw Dec (decision space) and Draw Obj (objective space) to observe the population’s evolution in real-time. Upon completion of the run, the final performance metrics are calculated, and their convergence behavior (as shown in Fig. 1 (e)) is displayed in a dedicated results panel on the right side of the interface.
The Test Module serves as an essential preliminary step for researchers to understand the nuances of specific MTO problems and the behavior of algorithms before embarking on large-scale experiments.